OpenStudy (anonymous):

–2(6 – 20n) = 20(2n – 3) A. 40 B. –36 C. no solution D. all real numbers

OpenStudy (twopointinfinity):

lol its not D

OpenStudy (anonymous):

expand

OpenStudy (twopointinfinity):

wait are we trying to find n?

OpenStudy (anonymous):

Duck knows the answer @TwoPointInfinity why bother helping duck out

OpenStudy (anonymous):

-2*6+2*20n=20*2n-3*20 -12+40n=40n-60 -12=-60 or 12=60 which is not true so no solution

OpenStudy (anonymous):

LOL ryan , at least they get medals so piss off @RyanL.

OpenStudy (anonymous):

At least im not an idiot who can't figure out a simple problem. Seriously do you really need help with this problem ? Is it really that hard for you?

OpenStudy (anonymous):

no , you got it right the first time ya dunce , i know the answer to every question i post , but i'm testing th knowlege of others , in the end whom ever helps the best and answers the question correctly gets a medal now once again piss off

OpenStudy (anonymous):

the*

OpenStudy (anonymous):

@RyanL.

OpenStudy (twopointinfinity):

You should block him so he doesn't appear on your posts.

OpenStudy (anonymous):

God are you really that stupid that think you are only posting question to test us out. That's not what this website is for.